OpenStudy (anonymous):

Write the nth term of the following sequence in terms of the first term of the sequence. 1, 8, 15, 22, . . .

OpenStudy (helder_edwin):

did you notice \[ \Large 8-1=7\] \[ \Large 15-8=7 \] \[ \Large 22-15=7 \]

OpenStudy (anonymous):

Tn=a+(n-1)d a=first term d=difference between every two terms n=number of the term

OpenStudy (anonymous):

wait it'll be 1+(n-1)7

OpenStudy (helder_edwin):

you were right with 7n-6

OpenStudy (anonymous):

It wont fit in the box they gave me

OpenStudy (helder_edwin):

so it would 1+4 times 7

OpenStudy (helder_edwin):

or 1+28

OpenStudy (anonymous):

??? no it needs to be to the nth term

OpenStudy (helder_edwin):

so it has to be 7n-6

OpenStudy (anonymous):

but that wont fit in the box

OpenStudy (helder_edwin):

i don't know! the answer is 1+7(n-1)

OpenStudy (anonymous):

ill try it lol...